OpenStudy (anonymous):

Medal - I need somebody to help show me how to break the example down correctly TY -5=s-3

OpenStudy (anonymous):

Can you give this a try first, it's the same as your last problem. @Nervonix Please don't give direct answers.

OpenStudy (anonymous):

c I got 2 only it's my signs

OpenStudy (eric_d):

Do it step-by-step

OpenStudy (anonymous):

s-3+=-5 +3 +3 -3+3=0 s=-5+3 s=-2 (-2)-3=-5 sorry about the first problem i forgot it was negative 5

OpenStudy (anonymous):

Ty I keep screwing up the neg with pos

OpenStudy (anonymous):

no Problem its tricky to get the hang off